Solve by any method you choose.
x-y+3z=13
4x+y+2z=17
3x+2y+2z=1

Answer by richwmiller(17219)

1,-1,3,13
4,1,2,17
3,2,2,1
I used determinants

x = 105/15 = 7
y = -135/15 = -9
z = -15/15 = -1